# NeuroPareto: Calibrated Acquisition for Costly Many-Goal Search in Vast Parameter Spaces

### v1

@@ 1 Introduction
Real-world engineering and scientific design problems often require simultaneously optimizing multiple conflicting objectives while each objective evaluation is costly in computation or experiment. Traditional population-based multi-objective search methods remain effective when many objective evaluations are affordable, but they lose practicality in constrained-budget settings. In response, surrogate-assisted strategies replace expensive evaluations with predictions from learned models to steer search. These approaches have matured into diverse families that emphasize either surrogate fidelity, acquisition design, or evolutionary mechanisms for preserving diversity and handling constraints. Representative work in surrogate-assisted multi-objective optimization and related algorithmic refinements illustrate both the promise and the limitations of these directions. (Pang et al., 2022; Chen et al., 2023; Zhou et al., 2024)
A parallel strand of research emphasizes learned components and classifier-guided screening as a way to reduce true evaluations. Classifier-based nondomination prediction and learned offspring generators have shown that many poor candidates can be filtered cheaply before committing costly evaluations. Hybrid pipelines that combine classifier screening with surrogate-managed infill extend the practical reach of surrogate assistance but expose two weaknesses. First, classifiers and surrogate models must provide well calibrated uncertainty so that screening does not systematically bias search away from true optima. Second, learned acquisition or history-aware policies must generalize from sparse optimization traces. Recent works on calibration, ensemble-style uncertainty approximations, and complexity-aware surrogate management address pieces of this picture while leaving open how to combine them effectively. (Yuan and Banzhaf, 2021; Durasov et al., 2021; Li et al., 2024b)
A further challenge arises from high dimensionality. Dimensionality reduction, random embedding, and representation learning relieve modeling burdens by focusing capacity on relevant subspaces. At the same time, scalable surrogate constructions and low-rank approximations make it feasible to reason about uncertainty under larger decision spaces. Despite these advances, co-design between classifier screening, uncertainty-aware surrogates, and adaptive acquisition remains uncommon. Methods that improve a single component in isolation often fail to capture the cross-component interactions that determine overall performance when evaluation budgets are severely limited. (Shmakov et al., 2023; Li et al., 2025; Liu et al., 2025) Although named NeuroPareto, the proposed method is fundamentally a hybrid Bayesian optimization algorithm; the suffix ‘EA’ emphasizes the use of rank-based population operators rather than indicating a traditional evolutionary algorithm.
@@ 2.1 Multi-objective Bayesian Optimization
Multi-objective Bayesian optimization (MOBO) offers a principled approach for expensive black-box problems with competing objectives. Classic hypervolume-aware and information-theoretic acquisition rules perform well in low dimensions but face scalability challenges (Daulton et al., 2021; Pang et al., 2022). Recent work improves EHVI and its multi-point variants for better tractability and stability (Deng et al., 2025), while generative and diffusion-based Pareto set models capture complex distributions under tight budgets (Li et al., 2025). For high-dimensional or medium-scale problems, hybrid surrogates and inverse-distance or radial-basis proxies trade fidelity for tractability (Li et al., 2024a; Liu et al., 2025).
@@ 2.2 Learned acquisition, meta-learning and surrogate policies
Recent work replaces hand-designed acquisition heuristics with learned policies that generalize across tasks. Meta-BO and task-conditioned surrogate frameworks, including transformer-based and amortized learners, jointly learn models and acquisition strategies to improve sample efficiency on related task families (Shmakov et al., 2023; Buathong et al., 2023). Structure-aware approaches that exploit partial evaluations or cost models further reduce evaluation expense (Song et al., 2022). Learned acquisition is promising for multi-objective settings but requires robust uncertainty estimates and compact feature representations to remain effective in high-dimensional spaces.
@@ 2.3 Uncertainty quantification and calibration
Accurate uncertainty quantification underpins effective exploration. Ensembles and Bayesian approximations remain the standard tools for decomposing epistemic and aleatoric uncertainty (Gal and Ghahramani, 2016; Abulawi et al., 2025). Practical methods aim to reduce the cost of ensemble-like behaviours while preserving calibration: Masksembles interpolates between MC-Dropout and deep ensembles, improving reliability at modest overhead (Durasov et al., 2021). Recent proposals focus on improving MC-Dropout calibration and integrating optimizer-based tuning for better uncertainty estimates in deep models (Asgharnezhad et al., 2025). In dynamic or parametric surrogate settings, explicitly modelling environment or observable parameters helps separate reducible from irreducible uncertainty and improves Pareto-front tracking under change (Temmerman et al., 2025).
@@ 2.4 Neural–GP hybrids and scalable surrogate design
Neural–GP hybrids and deep Gaussian process families aim to combine neural expressivity with Gaussian-process uncertainty quantification. Advances in amortized variational inference, sparse inducing schemes and randomized-feature approximations permit significantly larger training sets and deeper latent stacks while keeping uncertainty estimates meaningful (Meng and Zhang, 2024; Rochussen and Fortuin, 2025). For expensive many-variable tasks, two-level model management and low-rank factorization strategies have been proposed to control surrogate costs without discarding uncertainty-aware decision rules (Liu et al., 2025; Li et al., 2024b).
@@ 2.5 Classifier-assisted selection and learning-aided evolutionary methods
Classifier-guided selection, dominance prediction, and learned offspring generation reduce expensive evaluations by screening or biasing the evolutionary pipeline. Dominance predictors, rank-conditioned generation, and local surrogate infill accelerate convergence and preserve diversity when evaluations are scarce (Yuan and Banzhaf, 2021; Chen et al., 2023; Guo et al., 2024). Surrogate-assisted neuroevolution and linear-genetic-programming-based surrogate schemes show that surrogate paradigms extend beyond canonical EMO benchmarks and reduce computational cost (Stapleton et al., 2024). Feature-selection and permutation-based multi-objective strategies highlight the need for specialized high-dimensional operators (Espinosa et al., 2025). Algorithms for dynamic or constrained landscapes demonstrate benefits of multiple populations and explicit infeasible-region handling when feasible sets are irregular (Jiang et al., 2025).
@@ 2.6 Positioning of NeuroPareto
NeuroPareto is positioned at the intersection of classifier-assisted selection, uncertainty-aware surrogate modeling, and learned acquisition. It integrates calibrated rank prediction (to cheaply guide candidate generation), expressive Deep GP surrogates (to decompose predictive uncertainty), and a history-aware acquisition learner (to adapt evaluation priorities based on observed hypervolume gains). By combining these elements and adopting complexity-aware approximations, NeuroPareto aims to improve robustness and scalability on high-dimensional, expensive multi-objective tasks compared with methods that optimize individual components in isolation (Chen et al., 2023; Zhou et al., 2024; Li et al., 2025).
@@ 3.4.2 Temperature calibration and adaptive MC dropout
where $u_{\mathrm{ep}}^{\mathrm{clf}}(\mathbf{x})$ is nonnegative; this term corresponds to the information gain between the predictive label and the Dropout-induced posterior over model parameters, and thus quantifies epistemic uncertainty under the variational approximation(Gal and Ghahramani, 2016). See Appendix 14 for a detailed derivation and empirical validation. The classifier is evaluated in batched mode on large candidate pools and the pairs $\big(\bar{\mathbf{p}}(\mathbf{x}),u_{\mathrm{ep}}^{\mathrm{clf}}(\mathbf{x})\big)$ are cached to avoid repeated forward passes during screening and selection.
@@ 4.1 Experimental Configuration
Comparative Algorithms: We evaluate against ten surrogate-assisted evolutionary algorithms and four established MOBO baselines. CPS-MOEA (Zhang et al., 2015) steers search via a Pareto-dominance classifier, while K-RVEA (Chugh et al., 2016) adapts reference vectors with Kriging surrogates. CSEA (Pan et al., 2018) fuses multiple classifiers for decision support, EDN-ARMOEA (Lin et al., 2021) embeds neural dropout for uncertainty, and MCEA/D (Sonoda and Nakata, 2022) decomposes the problem with an ensemble of classifiers. CLMEA (Chen et al., 2023) refines local models through rank-based learning. In addition, GP-EI (Jones et al., 1998), RF-EI (Hutter et al., 2011), GP-HV (Daulton et al., 2020) and CL-EGO Zhang et al. (2015) serve as gold-standard Bayesian optimisation baselines.

### v2

@@ 1 Introduction
Real-world engineering and scientific design problems often require simultaneously optimizing multiple conflicting objectives while each objective evaluation is costly in computation or experiment. Traditional population-based multi-objective search methods remain effective when many objective evaluations are affordable, but they lose practicality in constrained-budget settings. In response, surrogate-assisted strategies replace expensive evaluations with predictions from learned models to steer search. These approaches have matured into diverse families that emphasize either surrogate fidelity, acquisition design, or evolutionary mechanisms for preserving diversity and handling constraints. Representative work in surrogate-assisted multi-objective optimization and related algorithmic refinements illustrate both the promise and the limitations of these directions. [26, 4, 35]
A parallel strand of research emphasizes learned components and classifier-guided screening as a way to reduce true evaluations. Classifier-based nondomination prediction and learned offspring generators have shown that many poor candidates can be filtered cheaply before committing costly evaluations. Hybrid pipelines that combine classifier screening with surrogate-managed infill extend the practical reach of surrogate assistance but expose two weaknesses. First, classifiers and surrogate models must provide well calibrated uncertainty so that screening does not systematically bias search away from true optima. Second, learned acquisition or history-aware policies must generalize from sparse optimization traces. Recent works on calibration, ensemble-style uncertainty approximations, and complexity-aware surrogate management address pieces of this picture while leaving open how to combine them effectively. [33, 11, 20]
A further challenge arises from high dimensionality. Dimensionality reduction, random embedding, and representation learning relieve modeling burdens by focusing capacity on relevant subspaces. At the same time, scalable surrogate constructions and low-rank approximations make it feasible to reason about uncertainty under larger decision spaces. Despite these advances, co-design between classifier screening, uncertainty-aware surrogates, and adaptive acquisition remains uncommon. Methods that improve a single component in isolation often fail to capture the cross-component interactions that determine overall performance when evaluation budgets are severely limited. [28, 18, 22] Although named NeuroPareto, the proposed method is fundamentally a hybrid Bayesian optimization algorithm; the suffix ‘EA’ emphasizes the use of rank-based population operators rather than indicating a traditional evolutionary algorithm.
@@ 2.1 Multi-objective Bayesian Optimization
Multi-objective Bayesian optimization (MOBO) offers a principled approach for expensive black-box problems with competing objectives. Classic hypervolume-aware and information-theoretic acquisition rules perform well in low dimensions but face scalability challenges [7, 26]. Recent work improves EHVI and its multi-point variants for better tractability and stability [9], while generative and diffusion-based Pareto set models capture complex distributions under tight budgets [18]. For high-dimensional or medium-scale problems, hybrid surrogates and inverse-distance or radial-basis proxies trade fidelity for tractability [19, 22].
@@ 2.2 Learned acquisition, meta-learning and surrogate policies
Recent work replaces hand-designed acquisition heuristics with learned policies that generalize across tasks. Meta-BO and task-conditioned surrogate frameworks, including transformer-based and amortized learners, jointly learn models and acquisition strategies to improve sample efficiency on related task families [28, 3]. Structure-aware approaches that exploit partial evaluations or cost models further reduce evaluation expense [29]. Learned acquisition is promising for multi-objective settings but requires robust uncertainty estimates and compact feature representations to remain effective in high-dimensional spaces.
@@ 2.3 Uncertainty quantification and calibration
Accurate uncertainty quantification underpins effective exploration. Ensembles and Bayesian approximations remain the standard tools for decomposing epistemic and aleatoric uncertainty [13, 1]. Practical methods aim to reduce the cost of ensemble-like behaviours while preserving calibration: Masksembles interpolates between MC-Dropout and deep ensembles, improving reliability at modest overhead [11]. Recent proposals focus on improving MC-Dropout calibration and integrating optimizer-based tuning for better uncertainty estimates in deep models [2]. In dynamic or parametric surrogate settings, explicitly modelling environment or observable parameters helps separate reducible from irreducible uncertainty and improves Pareto-front tracking under change [32].
@@ 2.4 Neural–GP hybrids and scalable surrogate design
Neural–GP hybrids and deep Gaussian process families aim to combine neural expressivity with Gaussian-process uncertainty quantification. Advances in amortized variational inference, sparse inducing schemes and randomized-feature approximations permit significantly larger training sets and deeper latent stacks while keeping uncertainty estimates meaningful [24, 27]. For expensive many-variable tasks, two-level model management and low-rank factorization strategies have been proposed to control surrogate costs without discarding uncertainty-aware decision rules [22, 20].
@@ 2.5 Classifier-assisted selection and learning-aided evolutionary methods
Classifier-guided selection, dominance prediction, and learned offspring generation reduce expensive evaluations by screening or biasing the evolutionary pipeline. Dominance predictors, rank-conditioned generation, and local surrogate infill accelerate convergence and preserve diversity when evaluations are scarce [33, 4, 14]. Surrogate-assisted neuroevolution and linear-genetic-programming-based surrogate schemes show that surrogate paradigms extend beyond canonical EMO benchmarks and reduce computational cost [31]. Feature-selection and permutation-based multi-objective strategies highlight the need for specialized high-dimensional operators [12]. Algorithms for dynamic or constrained landscapes demonstrate benefits of multiple populations and explicit infeasible-region handling when feasible sets are irregular [16].
@@ 2.6 Positioning of NeuroPareto
NeuroPareto is positioned at the intersection of classifier-assisted selection, uncertainty-aware surrogate modeling, and learned acquisition. It integrates calibrated rank prediction (to cheaply guide candidate generation), expressive Deep GP surrogates (to decompose predictive uncertainty), and a history-aware acquisition learner (to adapt evaluation priorities based on observed hypervolume gains). By combining these elements and adopting complexity-aware approximations, NeuroPareto aims to improve robustness and scalability on high-dimensional, expensive multi-objective tasks compared with methods that optimize individual components in isolation [4, 35, 18].
@@ 3.4.2 Temperature calibration and adaptive MC dropout
where $u_{\mathrm{ep}}^{\mathrm{clf}}(\mathbf{x})$ is nonnegative; this term corresponds to the information gain between the predictive label and the Dropout-induced posterior over model parameters, and thus quantifies epistemic uncertainty under the variational approximation[13]. See Appendix 14 for a detailed derivation and empirical validation. The classifier is evaluated in batched mode on large candidate pools and the pairs $\big(\bar{\mathbf{p}}(\mathbf{x}),u_{\mathrm{ep}}^{\mathrm{clf}}(\mathbf{x})\big)$ are cached to avoid repeated forward passes during screening and selection.
@@ 4.1 Experimental Configuration
Comparative Algorithms: We evaluate against ten surrogate-assisted evolutionary algorithms and four established MOBO baselines. CPS-MOEA [34] steers search via a Pareto-dominance classifier, while K-RVEA [5] adapts reference vectors with Kriging surrogates. CSEA [25] fuses multiple classifiers for decision support, EDN-ARMOEA [21] embeds neural dropout for uncertainty, and MCEA/D [30] decomposes the problem with an ensemble of classifiers. CLMEA [4] refines local models through rank-based learning. In addition, GP-EI [17], RF-EI [15], GP-HV [6] and CL-EGO [34] serve as gold-standard Bayesian optimisation baselines.